# You Only Learn One Query: Learning Unified Human Query for Single-Stage Multi-Person Multi-Task Human-Centric Perception

### 09 Dec 2023 v1

@@ 1 Introduction
Human-centric visual perception (e.g. pedestrian detection, pose estimation [^1] In this paper, we use “keypoint localization” and “pose estimation” interchangeably., human segmentation and human attribute recognition have attracted increasing research attention owing to their widespread industrial applications such as sports analysis, virtual reality, and augmented reality.
@@
Figure 1: Multi-person human-centric perception tasks can be categorized into 4 groups: classification, detection, segmentation and keypoint localization.
@@
The task of single-stage multi-person multi-task human-centric perception (HCP) has not been fully exploited in the literature due to the absence of a representative benchmark dataset. Consequently, previous studies [12] resorted to training models on various datasets for each HCP task, which can introduce certain limitations. Firstly, there is inherent scale variance across different datasets. For example, human detection datasets [46] consist of scene images with multiple interacting people, while attribute recognition datasets [52] typically contain images with a single cropped person. This hampers the development of single-stage multi-task algorithms that can comprehensively address various HCP tasks as a unified problem. Secondly, single-task datasets are often designed for specific application scenarios, resulting in strong dataset biases across different datasets. For example, some datasets [27] are captured in controlled lab environments, while some [52] are captured from a surveillance viewpoint. Naively training models on a combination of these datasets inevitably introduces dataset biases and hinders performance in real-world, unconstrained scenarios. Although there are separate benchmarks for individual HCP tasks, a comprehensive benchmark to simultaneously evaluate multiple HCP tasks is still lacking. To address this problem, we introduce a large-scale benchmark dataset called COCO-UniHuman, specifically designed for unified human-centric perceptions. As shown in Figure 1, most popular HCP tasks [^2] In this work, we focus on 2D human-centric perception tasks. can be grouped into four fundamental categories: classification, detection, segmentation, and keypoint localization. The COCO-UniHuman dataset extends COCO dataset by extensively annotating gender and age labels for each person instance. It encompasses all these four categories, covering 6 diverse HCP tasks (marked with check marks in Figure 1).
@@
Prior works on multi-person multi-task HCP have predominantly employed a multi-stage approach. These approaches typically involve employing a human detector to detect human instances, followed by task-specific models for each individual human perception task such as pose estimation and instance segmentation. However, these approaches exhibit three significant drawbacks. Firstly, they suffer from the issue of early commitment: the performance of the whole pipeline highly relies on body detection, and there is no recourse to recovery if the body detector fails. Secondly, the run-time is proportional to the number of individuals present in the image, making them computationally expensive for real-time applications. In contrast, single-stage methods estimate all required properties for human-centric analysis in a single pass, resulting in improved efficiency. Thirdly, these approaches overlook the potential inter-task synergy. Different HCP tasks are highly correlated as they share a common understanding of human body structure. In this work, we develop a simple, straightforward and versatile baseline framework, called HQNet, for single-stage multi-task HCP. It unifies various distinct human-centric tasks, including pedetrian detection, human segmentation, human pose estimation, and human attribute analysis (specifically gender and age).
@@
Different HCP tasks have their own relevant features of diverse granularity to focus on. For instance, pedestrian detection emphasizes global semantic features; attribute recognition necessitates both global and local semantic cues; person segmentation relies on fine-grained semantic features; and pose estimation require fine-grained semantic and localization information. In this paper, we propose to learn unified all-in-one query representations, termed Human Query, to encode instance-specific features of diverse granularity from multiple perspectives. Our work is inspired by DETR-based methods [8, 96, 51, 40, 106], which employ learnable query embeddings to represent objects and infer the relations of the objects and the image features. This study expands upon these works by learning versatile instance-level query representations for general human-centric perceptions. Despite the absence of intricate model designs, HQNet achieves state-of-the-art results across various benchmarks. Furthermore, we highlight several noteworthy characteristics of HQNet, including its flexibility, scalability, and transferability. (1) Flexibility: HQNet can readily integrate with diverse backbone networks, such as ResNet [23], Swin [54] and ViT [17]. (2) Scalibility: the weight-sharing backbone, transformer encoder, and decoder in HQNet enables seamless integration with multiple tasks, with minimal overhead from each task-specific head, thus demonstrating remarkable scalability. (3) Transferability: Experiments demonstrate strong transferability of the learned Human Query to novel HCP tasks, such as face detection and multi-object tracking. It should be noted that this work does not claim the algorithmic superiority but rather establishes a solid baseline with superior performance in human-centric perception.
@@
Our work makes the following key contributions: (1) We introduce the COCO-UniHuman benchmark, a large-scale dataset that comprehensively covers all representative HCP tasks, i.e. classification (gender and age estimation), detection (body and face detection), segmentation, and keypoint localization. (2) We develop a simple yet effective baseline called HQNet, unifying multiple distinctive HCP tasks in a single-stage multi-task manner. The key idea is to learn unified all-in-one query representations, termed Human Query, which encode instance-specific features of diverse granularity from various perspectives. Our approach achieves state-of-the-art results on different HCP tasks, demonstrating the strong representation capability of the learnt Human Query. Furthermore, experiments show the strong transferability of the learned Human Query to novel HCP tasks, such as face detection and multi-object tracking. We hope our work can shed light on future research on developing single-stage multi-person multi-task HCP algorithms.
@@
Dataset # Img # Inst # ID Crop Body Box Face Box Body Kpt Body Mask Gender Age
@@
2 Related Works
2.1 Human-centric perception tasks and datasets
Human-centric perception (HCP) tasks including pose estimation [87, 7, 29, 81, 78, 91], segmentation [24, 69, 22], and attribute recognition [38, 28, 68], have been extensively studied in computer vision. Approaches to multi-person HCP can be categorized into top-down, bottom-up, and single-stage methods. Top-down methods follow a detect-then-analyze approach. They first localize human instances, and then perform single person analysis. Top-down approaches can be divided into two types: those using separate pre-trained detectors and task-specific perception models [69], and those jointly learning detection and perception modules [24]. Bottom-up methods learn instance-agnostic keypoints/masks and cluster them using integer linear programming [26, 36, 30], heuristic greedy parsing [6, 65], embedding clustering [60, 37, 31], or learnable clustering [32]. Single-stage methods directly predict keypoints or masks for each individual, with different representations for pose estimation (coordinate-based [76, 63, 83, 93], heatmap-based [79, 71], or hybrid [105, 57, 20]) and segmentation (contour-based [88] or mask-based [5]). While most existing approaches focus on individual HCP tasks, we aim to unify human-centric perception by learning a single model that handles multiple tasks simultaneously, enabling a comprehensive understanding of humans.
As shown in Table 1, there are many task-specific datasets separately annotated for different HCP tasks, including pedestrian detection [16, 99, 70], pose estimation [3, 4, 85, 18], human segmentation [22, 42], and human attribute recognition [102, 52]. Datasets for multiple HCP tasks also exist. Widely used COCO [46, 33, 90] offers thorough annotations: body box, keypoints, and segmentation mask. Our COCO-UniHuman dataset further extends COCO featuring extensive gender and age annotations.
@@ 2.2 Unified methods for HCP
General network architecture for different HCP tasks. Some works design general network backbones, including CNN-based [80] and Transformer-based backbones [95]. Others unify HCP tasks with novel perception heads, such as UniHead [45]. Unlike these methods, which employ separate task-specific models, we consolidate diverse HCP tasks within a single network. Pre-training on HCP tasks. There are works [25, 10, 74] on pre-training on diverse human-centric tasks with large-scale data. UniHCP [12] presents a unified vision transformer model to perform multitask pre-training at scale. It employs task-specific queries for attending to relevant features, but tackles one task at a time. Unlike ours, our approach simultaneously solves multiple HCP tasks in a single forward pass. Our approach contrasts with these pre-training based methods by avoiding pre-training, minimizing fine-tuning, and circumventing resource-intensive multi-dataset training. Unlike them, we handle multiple HCP tasks concurrently in a single-stage, multi-task manner, diverging from their single-person focus. Co-learning on HCP tasks. Many works have investigated the correlations between pairs of HCP tasks [97, 75, 50, 61, 62]. We propose a single-stage model that learns a general unified representation to handle all representative HCP tasks simultaneously.
@@ 3 COCO-UniHuman Dataset
COCO-UniHuman is the first large-scale dataset, which provides annotations for all four representative HCP tasks in multi-person scenarios. Building upon COCO [46, 33], we have enriched the annotations by including gender and age information for each individual.
Uniqueness. The newly introduced dataset possesses several noteworthy properties in comparison to existing HCP datasets. (1) Comprehensiveness: This is the first large-scale multi-person HCP dataset that encompasses all four basic HCP tasks, i.e. classification, detection, segmentation, and keypoint localization in multi-person scenarios. It facilitates the development and evaluation of single-stage multi-person multi-task HCP algorithms. (2) Large scale and high diversity: With over 200,000 images and 273,000 identities, this dataset exhibits significant variations in terms of lighting conditions, image resolutions, human poses, and indoor/outdoor environments. (3) Multi-person attribute recognition: Unlike most existing human attribute recognition datasets that solely provide single-person center cropped images, our proposed dataset offers a valuable benchmark for multi-person attribute recognition in challenging scenarios. (4) Body-based apparent age estimation: While previous research has primarily focused on predicting a person’s age based on facial images, our dataset emphasizes the utilization of richer visual cues derived from whole-body images. Incorporating body-based visual cues such as skin elasticity, body posture, and body height proves beneficial for estimating a person’s age, particularly in situations where the facial image lacks clarity (e.g. captured from a distance). Notably, existing large-scale pedestrian attribute datasets [14] typically only offer coarse age group annotations, while facial attribute datasets [2] often provide fine-grained apparent or real age annotations. Our proposed dataset bridges this gap and serves as the pioneering large-scale dataset for body-based apparent age estimation in the wild. (5) Enhanced human representation: The extended attribute labels provide additional descriptive information about individuals beyond the existing labels. By leveraging these attribute labels, models can learn improved representations of humans, consequently enhancing the performance of other HCP tasks. Furthermore, the inclusion of gender and age labels can be valuable in downstream applications (e.g. gender/age-specific SMPL[55] model selection).
@@ 3.1 Data Annotation
To ensure accurate annotations, we employ trained annotators to manually label the gender and apparent age for each human instance in the dataset. We discard images full of non-human objects, and exclude all Small category persons that are hardly attribute-recognizable. Gender annotation. For each valid human instance, we adopt a body-based annotation approach. Using the provided human bounding boxes, we crop the body images and request annotators to label the gender. To maintain data quality, we conduct quality inspections and manual corrections throughout the labeling process. Age annotation. To enhance the quality of annotation, we employ a two-stage strategy based on body-based annotation. Similar to gender annotation, age annotation is also performed on cropped body images. We implement a coarse-to-fine two-stage annotation strategy, considering age group annotation to be comparatively easier than apparent age annotation [2]. In the first stage, age groups are annotated. Following [44], we divide the age ranges into six groups, i.e. “baby”, “kid”, “teen”, “young”, “middle aged”, and “elderly”. For each cropped person image, we request a group of 10 annotators to independently and repeatedly label the age groups (6-category classification task). We take the mode of the 10 votes as the ground-truth age group. In the second stage, the apparent age is annotated. Given the age group as a prior, a group of 10 annotators independently and repeatedly annotate the apparent age. Consequently, we obtain 10 votes for each human instance. We remove the outliers and take the average as the final ground-truth apparent age. As a summary, the dataset contains over 1M apparent votes. Experiments validate the effectiveness of the body-based annotation strategy and the two-stage annotation strategy (see Sec. A2).
@@ 4.1 Overview
This study endeavors to develop a single-stage framework that effectively supports a wide range of human-centric perception (HCP) tasks. The key is to learn a comprehensive human representation, which can be universally employed across various HCP tasks. To achieve this, we employ a query-based methodology and investigate the feasibility of representing each human instance as a single shared query.
Our framework is characterized by simplicity, flexibility, and scalability. Unlike previous task-specific HCP models that may incorporate specialized designs tailored to specific tasks (e.g. “mask-enhanced anchor box initialization” in Mask DINO [41]), our approach aims to handle various human-centric analysis tasks in a unified manner. To maximize knowledge sharing among various HCP tasks, we attempt to share most weights across different HCP tasks.
@@ 4.2 Task-shared Transformer Decoder
Queries in DETR-like models are formed by two parts: positional queries and content queries. Each positional query is formulated as a 4D anchor box, encoding the center x-y coordinates, width and height of the box, respectively. Our content query, denoted as Human Query, encapsulates various features (local and global appearance features, as well as coarse- and fine-grained localization features) specific to each instance.
Contrastive DeNoising (CDN). To enhance training stability and acceleration, we employ Contrastive DeNoising (CDN) as introduced in DINO [96]. Notably, we observe that incorporating auxiliary DeNoise losses for other tasks (e.g. segmentation and pose) does not yield significant improvements. Consequently, we only apply DN losses for human detection.
HumanQuery-Instance Matching. To ensure consistent and unique predictions for each ground-truth instance across all HCP tasks, including classification (Cls.), detection (Det.), keypoint (Kpt.), and segmentation (Seg.), we employ HumanQuery-Instance (HQ-Ins) Matching. $\lambda_{cls}L_{cls}+\lambda_{det}L_{det}+\lambda_{seg}L_{seg}+\lambda_{kpt}L_{kpt},$ where $\lambda$ are the corresponding loss weights (see Sec. A4 for details).

### 14 Mar 2024 v2

@@ Abstract
Keywords:
Human-Centric Perception Unified Vision Model
@@ 1 Introduction
Human-centric visual perception (e.g. pedestrian detection, 2D keypoint estimation, 3D mesh recovery, human segmentation and attribute recognition) have attracted increasing research attention owing to their widespread industrial applications such as sports analysis, virtual reality, and augmented reality.
@@
Figure 1: Multi-person human-centric perception tasks can be categorized into 4 groups: classification, detection, segmentation and pose estimation.
@@
The task of single-stage multi-person multi-task human-centric perception (HCP) has not been fully exploited in the literature due to the absence of a representative benchmark dataset. Consequently, previous studies [11] resorted to training models on various datasets for each HCP task, which can introduce certain limitations. Firstly, there is inherent scale variance across different datasets. For example, human detection datasets [45] consist of scene images with multiple interacting people, while attribute recognition datasets [49] typically contain images with a single cropped person. This hampers the development of single-stage multi-task algorithms that can comprehensively address various HCP tasks as a unified problem. Secondly, single-task datasets are often designed for specific application scenarios, resulting in strong dataset biases across different datasets. For example, some datasets [24] are captured in controlled lab environments, while some [49] are captured from a surveillance viewpoint. Naively training models on a combination of these datasets inevitably introduces dataset biases and hinders performance in real-world, unconstrained scenarios. Although there are separate benchmarks for individual HCP tasks, a comprehensive benchmark to simultaneously evaluate multiple HCP tasks is still lacking. To address this problem, we introduce a large-scale benchmark dataset called COCO-UniHuman, specifically designed for unified human-centric perceptions. As shown in Figure 1, most popular HCP tasks can be grouped into four fundamental categories: classification, detection, segmentation, and pose estimation. The COCO-UniHuman dataset extends COCO dataset by extensively annotating gender and age labels for each person instance. It encompasses all these four categories, covering 7 diverse HCP tasks (marked with check marks in Figure 1).
@@
Prior works on multi-person multi-task HCP have predominantly employed a multi-stage approach. These approaches typically involve employing a human detector to detect human instances, followed by task-specific models for each individual human perception task such as keypoint estimation and instance segmentation. However, these approaches exhibit three significant drawbacks. Firstly, they suffer from the issue of early commitment: the performance of the whole pipeline highly relies on body detection, and there is no recourse to recovery if the body detector fails. Secondly, the run-time is proportional to the number of individuals present in the image, making them computationally expensive for real-time applications. In contrast, single-stage methods estimate all required properties for human-centric analysis in a single pass, resulting in improved efficiency. Thirdly, these approaches overlook the potential inter-task synergy. Different HCP tasks are highly correlated as they share a common understanding of human body structure. In this work, we develop a simple, straightforward and versatile baseline framework, called HQNet, for single-stage multi-task HCP. It unifies various distinct human-centric tasks, including pedetrian detection, human segmentation, 2D human keypoint estimation, 3D human mesh recovery, and human attribute analysis (specifically gender and age).
@@
Different HCP tasks have their own relevant features of diverse granularity to focus on. For instance, pedestrian detection emphasizes global semantic features; attribute recognition necessitates both global and local semantic cues; person segmentation relies on fine-grained semantic features; and pose estimation require fine-grained semantic and localization information. In this paper, we propose to learn unified all-in-one query representations, termed Human Query, to encode instance-specific features of diverse granularity from multiple perspectives. Our work is inspired by DETR-based methods [7, 85, 48, 35, 95], which employ learnable query embeddings to represent objects and infer the relations of the objects and the image features. This study expands upon these works by learning versatile instance-level query representations for general human-centric perceptions. In addition, we design HumanQuery-Instance Matching (HQ-Ins Matching) and Gender-aided human Model Selection (GaMS) mechanisms to further exploit the interactions among different HCP tasks and enhance the performance of multi-task HCP.
We highlight several noteworthy characteristics of HQNet. (1) Flexibility: HQNet can readily integrate with diverse backbone networks, such as ResNet [21], Swin [50] and ViT [16]. (2) Scalibility: the weight-sharing backbone, transformer encoder, and decoder in HQNet enables seamless integration with multiple tasks, with minimal overhead from each task-specific head, thus demonstrating remarkable scalability. (3) Transferability: Experiments demonstrate strong transferability of the learned Human Query to novel HCP tasks, such as face detection and multi-object tracking.
@@
Our work makes the following key contributions: (1) We introduce the COCO-UniHuman benchmark, a large-scale dataset that comprehensively covers all representative HCP tasks, i.e. classification (gender and age estimation), detection (body and face detection), segmentation, and pose estimation (2D keypoint and 3D mesh recovery). (2) We develop a simple yet effective baseline called HQNet, unifying multiple distinctive HCP tasks in a single-stage multi-task manner. The key idea is to learn unified all-in-one query representations, termed Human Query, which encode instance-specific features of diverse granularity from various perspectives. Additionally, we design HumanQuery-Instance (HQ-Ins) Matching and Gender-aided human Model Selection (GaMS) mechanisms to improve the performance of multi-task HCP. (3) Our approach achieves state-of-the-art results on different HCP tasks, demonstrating the strong representation capability of the learnt Human Query. Furthermore, experiments show the strong transferability of the learned Human Query to novel HCP tasks, such as face detection and multi-object tracking. We hope our work can shed light on future research on developing single-stage multi-person multi-task HCP algorithms.
2 Related Works
2.1 Human-Centric Perception Tasks and Datasets
Human-centric perception (HCP) tasks including keypoint estimation, segmentation, and attribute recognition, have been extensively studied in computer vision. Approaches to multi-person HCP can be categorized into top-down, bottom-up, and single-stage methods. Top-down methods follow a detect-then-analyze approach. They first localize human instances, and then perform single person analysis. Top-down approaches can be divided into two types: those using separate pre-trained detectors and task-specific perception models [63, 30, 84, 42, 67, 74], and those jointly learning detection and perception modules [20]. Bottom-up methods learn instance-agnostic keypoints/masks and cluster them using integer linear programming [23, 32], heuristic greedy parsing [6, 59], embedding clustering [54, 33], or learnable clustering [26]. Single-stage methods directly predict keypoints or masks for each individual, with different representations for 2D keypoint estimation (coordinate-based [71, 56, 76, 81], heatmap-based [73, 66], or hybrid [94, 53, 18]), 3D mesh recovery [68, 41] and segmentation (contour-based [79] or mask-based [5]). While existing approaches focus on individual HCP tasks, we aim to unify HCP by learning a single model that handles multiple tasks simultaneously, enabling a comprehensive understanding of humans.
@@ 2.1 Human-Centric Perception Tasks and Datasets
Dataset # Img # Inst # ID Crop Body Box Face Box Body Kpt Body Mask Gender Age Mesh
@@
As shown in Table 1, there are many task-specific datasets separately annotated for different HCP tasks, including pedestrian detection [15, 88, 64], multi-human keypoint estimation [4, 3], multi-human segmentation [19, 37], and human attribute recognition [92, 49]. Datasets for multiple HCP tasks also exist. Widely used COCO [45, 27] offers thorough annotations: body box, keypoints, and segmentation mask. Our COCO-UniHuman dataset further extends COCO featuring extensive gender and age annotations.
@@ 2.2 Unified Methods for HCP
General network architecture for different HCP tasks. Some works design general network backbones, including CNN-based [74] and Transformer-based backbones [83]. Others unify HCP tasks with novel perception heads, such as UniHead [40]. Unlike these methods, which employ separate task-specific models, we consolidate diverse HCP tasks within a single network. Pre-training on HCP tasks. There are also works [22, 9, 69] on pre-training on diverse human-centric tasks with large-scale data. More recently, UniHCP [11] presents a unified vision transformer model to perform multitask pre-training at scale. It employs task-specific queries for attending to relevant features, but tackles one task at a time. Unlike ours, our approach simultaneously solves multiple HCP tasks in a single forward pass. Our approach contrasts with these pre-training based methods by avoiding pre-training, minimizing fine-tuning, and circumventing resource-intensive multi-dataset training. Unlike them, we handle multiple HCP tasks concurrently in a single-stage, multi-task manner, diverging from their single-person focus. Co-learning on HCP tasks. Many works have investigated the correlations between pairs of HCP tasks [86, 70, 46, 55, 57]. We propose a single-stage model that learns a general unified representation to handle all representative human-centric perception tasks simultaneously.
@@ 3 COCO-UniHuman Dataset
COCO-UniHuman is the first large-scale dataset, which provides annotations for all four representative HCP tasks in multi-person scenarios. Building upon MSCOCO’2017 [45, 27, 28] dataset, we have enriched the annotations by including gender, age, 3D body mesh information for each individual.
@@ 3.1 Data Annotation
To ensure accurate annotations, we employ trained annotators to manually label the gender and apparent age for each human instance in the dataset. We discard images full of non-human objects, and exclude all Small category persons that are hardly attribute-recognizable. Gender annotation. For each valid human instance, we adopt a body-based annotation approach. Using the provided human bounding boxes, we crop the body images and request annotators to label the gender. To maintain data quality, we conduct quality inspections and manual corrections throughout the labeling process. Age annotation. To enhance the quality of annotation, we employ a two-stage strategy based on body-based annotation. Similar to gender annotation, age annotation is also performed on cropped body images. We implement a coarse-to-fine two-stage annotation strategy, considering age group annotation to be comparatively easier than apparent age annotation [2]. In the first stage, age groups are annotated. Following [39], we divide the age ranges into six groups, i.e. “baby”, “kid”, “teen”, “young”, “middle aged”, and “elderly”. For each cropped person image, we request a group of 10 annotators to independently and repeatedly label the age groups (6-category classification task). We take the mode of the 10 votes as the ground-truth age group. In the second stage, the apparent age is annotated. Given the age group as a prior, a group of 10 annotators independently and repeatedly annotate the apparent age. Consequently, we obtain 10 votes for each human instance. We remove the outliers and take the average as the final ground-truth apparent age. As a summary, the dataset contains over 1M apparent votes. Experiments validate the effectiveness of the body-based annotation strategy (Sec. A2.1) and the two-stage annotation strategy (Sec. A2.2). Mesh Annotation. We follow [28] to apply Exemplar Fine-Tuning (EFT) method to generate 3D pseudo-ground truth SMPL parameters. To ensure data quality, we only retain instances with at least 12 keypoint annotations where all limbs are visible and filter out low-quality data manually.
3.2 Data Uniqueness
The newly introduced dataset possesses several noteworthy properties in comparison to existing HCP datasets. (1) Comprehensiveness: This is the first large-scale multi-person HCP dataset that encompasses all four basic HCP tasks, i.e. classification, detection, segmentation, keypoint localization, and 3D body mesh recovery in multi-person scenarios. It facilitates the development and evaluation of single-stage multi-person multi-task HCP algorithms. (2) Large scale and high diversity: With over 200,000 images and 273,000 identities, this dataset exhibits significant variations in terms of lighting conditions, image resolutions, human poses, and indoor/outdoor environments. (3) Multi-person attribute recognition: Unlike most existing human attribute recognition datasets that solely provide single-person center cropped images, our proposed dataset offers a valuable benchmark for multi-person attribute recognition in challenging scenarios. (4) Body-based apparent age estimation: While previous research has primarily focused on predicting a person’s age based on facial images, our dataset emphasizes the utilization of richer visual cues derived from whole-body images. Incorporating body-based visual cues such as skin elasticity, body posture, and body height proves beneficial for estimating a person’s age, particularly in situations where the facial image lacks clarity (e.g. captured from a distance). Notably, existing large-scale pedestrian attribute datasets [13] typically only offer coarse age group annotations, while facial attribute datasets [2] often provide fine-grained apparent or real age annotations. Our proposed dataset bridges this gap and serves as the pioneering large-scale dataset for body-based apparent age estimation in the wild. (5) Enhanced human representation: The extended human attribute labels and 3D human mesh information provide additional descriptive information about individuals beyond the existing labels. By leveraging these information, models can learn improved representations of humans, consequently enhancing the performance of other HCP tasks.
@@ 4.1 Overview
This study endeavors to develop a single-stage framework that supports a wide range of human-centric perception (HCP) tasks. The key is to learn a comprehensive human representation, which can be universally employed across various HCP tasks. To achieve this, we employ a query-based methodology and investigate the feasibility of representing each human instance as a single shared query. Our objective is to build a framework that is characterized by simplicity, flexibility, and scalability. Unlike previous task-specific HCP models that may incorporate specialized designs tailored to specific tasks (e.g. “mask-enhanced anchor box initialization” in Mask DINO [36]), our approach aims to handle various human-centric analysis tasks in a unified manner. To maximize knowledge sharing among various HCP tasks, we attempt to share most weights across different HCP tasks.
@@ 4.2 Task-Shared Transformer Decoder
Queries in DETR-like models are formed by two parts: positional queries and content queries. Each positional query is formulated as a 4D anchor box, encoding the center x-y coordinates, width and height of the box, respectively. Our content query, denoted as Human Query, encapsulates various features (local and global appearance features, as well as coarse- and fine-grained localization features) specific to each instance. To enhance training stability and acceleration, we employ Contrastive DeNoising (CDN) as introduced in DINO [85]. Notably, we observe that incorporating auxiliary DeNoise losses for other tasks (e.g. segmentation and pose) does not yield significant improvements. Consequently, we only apply DN losses for human detection.
HumanQuery-Instance Matching. To ensure consistent and unique predictions for each ground-truth instance across all HCP tasks, including classification (Cls.), detection (Det.), pose (Pose.), and segmentation (Seg.), we employ HumanQuery-Instance (HQ-Ins) Matching. $\lambda_{cls}L_{cls}+\lambda_{det}L_{det}+\lambda_{seg}L_{seg}+\lambda_{pose}L_{pose},$ where $\lambda$ are the corresponding loss weights. Further details can be found in Sec. A4.1.
4.3 Task-Specific Heads
To ensure scalability, we categorize HCP tasks into three groups and design specific implementation paradigms for each category. Coordinate prediction tasks (e.g. object detection and keypoint estimation) share common reference points with bounding box prediction and directly regress the normalized offsets of each point. Dense prediction tasks (e.g. instance segmentation and human parsing) follow the design of Mask DINO [36], which involves constructing a high-resolution pixel embedding map by integrating features from both the backbone and the Transformer encoder. By performing a dot-product operation between the content query embedding and the pixel embedding map, an instance-aware pixel embedding map is generated, facilitating pixel-level classification. Classification tasks (e.g. determining if an instance is human, gender and age estimation) directly map the Human Query to the classification prediction results, as the Human Query inherently encodes the positional information.
To minimize the overhead of incorporating new tasks, we employ lightweight task-specific heads comprising simple linear layers. The details of our four task-specific heads are as follows: Human detection head. A 3-layer multi-layer perceptron (MLP) with a hidden dimension of $d$ is utilized to predict the normalized center x-y coordinates, height, and width of the bounding box w.r.t. the input image. Additionally, a linear projection layer (FC) is employed to predict the class label (human or non-human). 2D keypoint estimation head. Following the coordinate prediction paradigm, the learned Human Query is fed into a pose regression head (MLP) to regress the relative pose offsets w.r.t. the shared reference points of the detection head. A confidence prediction head (FC) is used to predict confidence score of having visible keypoints. Following PETR [65], joint decoder layers are employed to refine body poses by leveraging structured relations between body keypoints. An auxiliary heatmap branch is used to aid training and discarded during testing. Human instance segmentation head. A 3-layer MLP is used to process the instance-aware pixel embedding map and output a one-channel mask, which is then upsampled to match the original input image size. Human attribute head. The gender estimation head and the age estimation head operate in parallel. Both heads consist of two-layer MLPs. Gender estimation involves binary classification, while age estimation is formulated as an 85-class ([1, 85]) classification with softmax expected value [62] estimation. 3D mesh recovery head. Two 3-layer MLPs with the same hidden dimension are used to predict the pose and shape parameters respectively. These parameters are then fed into a SMPL body model to generate the 3D body meshes.
4.3.1 Gender-aided human Model Selection (GaMS).
There are three versions of SMPL models: male, female and neutral. Previous works usually use the neutral model because of the lack of gender annotation. COCO-UniHuman has both gender and 3D mesh annotations. To improve the performance of 3D mesh, we employ Gender-aided human Model Selection (GaMS) which selects different SMPL models by gender labels during the training and the inference stage.
5 Experiments